A telemarketer makes successful calls 10% of the time. What is the probability the telemarketer makes exactly 5 unsuccessful calls before the third sale is made?
Answer #1

P[ successful call ] = 10% = 0.1

The population proportion of success is p = 0.1, also, 1 - p = 1 - 0.1 = 0.9, and the required number of successes is r = 3.

We need to compute P[ X = 8 ] ( 5 failure, 3 success total 8 trials )

Therefore, we get that

P[ X = 8 ] =

P[ X = 8 ] = 0.0124